OpenStudy (anonymous):

Solve the initial value problem. \[(x^2-\sin^2(y))dx+xsin(2y)dy = 0, y(1) = \frac{ \pi }{ 3 }\] Find x.

OpenStudy (kainui):

First off, can you show that this is exact or not?

OpenStudy (anonymous):

The equation is not exact.

OpenStudy (kainui):

Alright, so do you know how to make it exact with an integrating factor?

OpenStudy (anonymous):

\[M _{y} = -2sinycosy\] \[N_{x} = 2sinycosy\] These two are not exact so you can find \[\frac{ M_{y} - N_{x} }{ N } = \frac{ -2 }{ x }\]

OpenStudy (anonymous):

Integrating factor is \[x ^{-2}\]

OpenStudy (kainui):

Alright, cool. So let's keep this ball rolling, what's next?

OpenStudy (anonymous):

Multiplying by the integrating factor gives the equation \[(1-x^{-2}\sin^2y)dx + (2x^{-1}sinycosy)dy = 0\] The equations are now exact.

OpenStudy (kainui):

Alright, so what's next?

OpenStudy (anonymous):

The general solution should be \[\psi(x,y) = \int\limits_{}^{}(1-x^{-2}\sin^2y)dx = x+x^{-1}\sin^2y + c(y)\] Therefore: \[\psi _{y}' = 2x^{-1}sinycosy+c'(y) = N = 2x^{-1}sinycosy\] Meaning c'(y) = 0 so \[c(y) = c\] Which should give the general solution \[\psi(x,y) = x+x^{-1}\sin^2y=c\]

OpenStudy (anonymous):

Using \[y(1) = \frac{ \pi }{ 3 }\] \[1 + 1^{-1}\sin^2(\frac{ \pi }{ 3 }) = c\] Therefore c = 7/4 Which means \[x + x^{-1}\sin^2y = \frac{ 7 }{ 4 }\] Is this correct?

OpenStudy (kainui):

Looks correct to me.

OpenStudy (anonymous):

How do I solve x as a function of y using that solution?

OpenStudy (loser66):

just isolate y sin^2y = -x^2 +(7/4)x

OpenStudy (loser66):

take square root both sides and arcsin both sides to get y respect to x

OpenStudy (anonymous):

The question is looking for x with respect to y, not y respect to x.

OpenStudy (loser66):

If so, you have x^2 -7/4 x + sin^2 y =0 right? solve as a quadratic problem with a =1, b = -7/4 and c = sin^2 y you may have 2 roots x1, and x2 consider the condition to get the right answer, you may have both roots . Got me?

OpenStudy (anonymous):

@Loser66 is right, here are the roots \[ x= \frac{1}{8} \left(7-\sqrt{49-64 \sin ^2(y)}\right)\\x= \frac{1}{8} \left(7+\sqrt{49-64 \sin ^2(y)}\right) \]